OpenStudy (anonymous):

6. Use factoring to solve the quadratic equations below, if possible. a. 2x2 + 3x - 54 = 0

mathslover (mathslover):

\[\large{2x^2+3x-54=2x^2+12x-9x-54}\] right?

mathslover (mathslover):

\[\large{2x^2+12x-9x-54=0}\] \[\large{2x(x+6)-9(x+6)=0}\] \[\large{(2x-9)(x+6)=0}\] \[\large{\textbf{Hence, either 2x-9=0 or x+6=0}}\] \[\large{2x-9=0 \space 2x=9 ,\space x=\frac{9}{2}}\] \[\large{x+6=0 ; x=-6}\]

OpenStudy (anonymous):

thanks